A large turntable rotates about a fixed vertical a
Answer #1

2.5 rad/sec = 23.873241491063457 RPM
23.873241491063457 RPM = 1 rotation in 2.5 sec
ω = 2π/ T =2π/ 2.5 = 2.51327 radian /s
Initial angular momentum = 900*2.513
Final angular momentum [900 + 40 *2^2] ω2
Angular momentum is conserved.
[900 + 40 *2^2] ω2 = 900 ω1
ω2 = 900 *2.51327 / [900 + 40 *2^2]
ω2 = 2.134 radian /s